Circular No. 8892

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 17P/HOLMES B. Yang and D. Jewitt, University of Hawaii, report near-infrared (0.8-

to 4.2-μm) spectra of comet 17P taken with the 3.0-m NASA Infrared Telescope Facility by S. J. Bus on Oct. 27 and 28 UT. Two broad absorption bands were found centered at 2.02 μm (the depth being 10 percent of the continuum) and 3.03 μm (30-percent deep), respectively, in the reflectance spectra of 17P. These features, which appeared on both nights, are consistent with abundant water ice grains in the central coma. The overall slope of the infrared reflectance spectrum of this comet is blue, with a normalized reflectance gradient near −2.3 percent per 100 nm.

COMETS C/2007 N4–N8, C/2007 O3–O4, C/2007 Q4–Q7 (SOHO) Further to IAUC 8891, additional Kreutz-sungrazing comets have been

found on SOHO website images, their “discovery” observations tabulated below. All eleven objects were stellar in appearance in C3 images. C/2007 N4, of mag 5.5 in C3 images, was slightly diffuse with no tail in C2 images. C/2007 N5, C/2007 N7, C/2007 Q4, C/2007 Q5, and C/2007 Q7 were of mag ∼ 6 in C3 images; C2 images showed C/2007 N5 to be quite diffuse with a hint of faint tail, and C/2007 N7 to be slightly diffuse with a very faint short tail. C/2007 N6, C/2007 N8, and C/2007 O4 were faint (mag ∼ 7) in C3 images; however, C/2007 N6 appeared diffuse with no tail in C2 images. C/2007 O3 was tiny and very faint (mag 7.5). C/2007 Q6 was very small and of mag ∼ 6.5. C/2007 Q7 was possibly slightly elongated.

Comet  2007 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2007 N4 July 3.613 6^{h}37.^{m}2 +19^{o}41′ C3/2 HS 2007-U23
C/2007 N5  7.904 6 52.8 +19 38 C3/2 BZ 2007-U23
C/2007 N6  11.638 7 09.2 +20 06 C3/2 BZ 2007-U23
C/2007 N7  13.179 7 11.2 +18 40 C3/2 KB 2007-U24
C/2007 N8  13.346 7 14.6 +19 35 C3  BZ 2007-U24
C/2007 O3  18.529 7 35.4 +19 43 C3  HS 2007-U24
C/2007 O4  27.221 8 05.9 +17 05 C3  BZ 2007-U24
C/2007 Q4 Aug. 17.238 9 22.4 +12 01 C3  SY 2007-U34
C/2007 Q5  17.263 9 22.8 +12 22 C3  AK 2007-U34
C/2007 Q6  17.529 9 28.1 +12 18 C3  HS 2007-U34
C/2007 Q7  25.096 9 50.2 + 9 37 C3  HS 2007-U34
2007 November 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green